What are the odds against a horse winning if the probability of the horse winning is 15/22?
Answer by stanbon(75887)

What are the odds against a horse winning if
the probability of the horse winning is 15/22
then probability of the horse losing is 7/22
------
So, odds against winning = P(losing)/P(winning) = 7/15 = 7:15
=================================================================
Cheers,
Stan H.
=========